# FreeGraftor: Training-Free Cross-Image Feature Grafting for Subject-Driven Text-to-Image Generation

### v3

@@ Abstract.
Subject-driven image generation aims to synthesize novel scenes that faithfully preserve subject identity from reference images while adhering to textual guidance. However, existing methods struggle with a critical trade-off between fidelity and efficiency. Tuning-based approaches rely on time-consuming and resource-intensive, subject-specific optimization, while zero-shot methods often fail to maintain adequate subject consistency. In this work, we propose FreeGraftor, a training-free framework that addresses these limitations through cross-image feature grafting. Specifically, FreeGraftor leverages semantic matching and position-constrained attention fusion to transfer visual details from reference subjects to the generated images. Additionally, our framework introduces a novel noise initialization strategy to preserve the geometry priors of reference subjects, facilitating robust feature matching. Extensive qualitative and quantitative experiments demonstrate that our method enables precise subject identity transfer while maintaining text-aligned scene synthesis. Without requiring model fine-tuning or additional training, FreeGraftor significantly outperforms existing zero-shot and training-free approaches in both subject fidelity and text alignment. Furthermore, our framework can seamlessly extend to multi-subject generation, making it practical for real-world deployment. Our code is available at https://github.com/Nihukat/FreeGraftor.
subject-driven image generation, text-to-image synthesis, training-free framework
$c$s: Computing methodologies Image manipulation
1. Introduction
The democratization of text-to-image generation has ignited widespread interest in personalized content creation, where users seek to synthesize novel scenes containing specific visual concepts—be it a cherished pet, a unique artifact, or a fictional character. While diffusion models(Nichol et al., 2021; Saharia et al., 2022; Ramesh et al., 2022; Rombach et al., 2022; Podell et al., 2023) excel at generating diverse imagery from textual prompts, adapting them to preserve subject identity without incurring costly per-subject optimization remains a critical challenge.
Current subject-driven generation methods(Ruiz et al., 2023; Gal et al., 2022; Kumari et al., 2023; Li et al., 2023; Ye et al., 2023) face a critical dilemma between fidelity and efficiency. Tuning-based methods (e.g., DreamBooth(Ruiz et al., 2023)) achieve high subject consistency by fine-tuning the model on reference images. However, this subject-specific optimization is resource-intensive, time-consuming, prone to catastrophic forgetting, and struggles with multi-subject composition. Zero-shot methods (e.g., IP-Adapter(Ye et al., 2023)) circumvent fine-tuning by using auxiliary modules (e.g., image encoders or adapters) to extract and inject subject identity information into the generation process. While efficient, these methods often fail to retain fine visual details, particularly for novel concepts, as the limited capacity of the auxiliary modules fundamentally restricts high-fidelity reconstruction.
Figure 1. Cross-image semantic correspondence in FLUX.1. For each specified pixel in the reference image, we identify the most similar pixel in the target image within FLUX.1‘s feature space. These corresponding pixel pairs are then visually connected by colored lines.
Critically, pre-trained text-to-image models inherently possess powerful visual representation capabilities, as evidenced by their ability to synthesize intricate textures and structures directly from text. However, current approaches largely overlook this intrinsic potential: tuning-based methods shift the generation distribution via fine-tuning, while zero-shot methods delegate representation learning to additional modules. Instead, directly leveraging the base model’s own rich feature space for cross-image knowledge transfer presents a promising avenue for bridging the fidelity-efficiency gap. Recent advances in appearance transfer(Epstein et al., 2023; Alaluf et al., 2024; Mou et al., 2023, 2024; Go et al., 2024) support this direction, demonstrating that diffusion models can transplant visual attributes (e.g., style, texture) between images through semantic feature matching. For instance, Eye-for-an-eye (Go et al., 2024) successfully transfers localized appearances by establishing cross-image correspondences within the U-Net’s feature space.
Building upon this concept, we identify a key insight: such cross-image semantic correspondences are not limited to U-Net architectures but also exist robustly within the feature space of modern MM-DiT-based models (Esser et al., 2024) (e.g., FLUX.1 (Labs, 2024)). As demonstrated in Fig. 1, computing pixel-wise similarity (detailed in Sec. 3.2) reveals that semantically corresponding pixels exhibit high similarity in FLUX.1’s feature space.
Figure 2. Subject-driven generation results of our FreeGraftor, achieved without any training or tuning process. These results demonstrate the superior properties of FreeGraftor, including (a) pixel-level detail preservation, (b) flexible text-guided control, and (c) support for multiple reference subjects.
@@ 1. Introduction
Leveraging this insight, we propose FreeGraftor, a training-free framework that resolves the fidelity-efficiency dilemma in subject-driven generation via cross-image feature grafting. Our approach extracts visual features from reference images using rectified flow inversion (Rout et al., 2024; Wang et al., 2024b; Deng et al., 2024) and injects them into the generation process. Specifically, we introduce a novel cross-image feature grafting mechanism that transfers appearance details through semantic matching and position-constrained attention fusion in the feature space of text-to-image diffusion models. To bind reference patches to their corresponding generated patches, we implement a position embedding replication strategy, enabling semantically matched patches to share positional information during attention computation. Furthermore, we identify that proper noise initialization is crucial for robust semantic matching and structural preservation of reference subjects. Thus, we propose a structure-consistent initialization strategy: first constructing a collage containing the reference subject, then inverting it as the initialization noise for generation. Our method achieves pixel-level detail preservation (e.g., text, patterns) with fast inference and low computational overhead, requiring no additional training or optimization.
@@
Extensive experiments demonstrate that our method outperforms existing approaches in both subject consistency and text alignment while maintaining competitive efficiency. When combining multiple subjects, FreeGraftor preserves individual identities and harmoniously integrates all subjects into a coherent scene without layout conflicts or attribute confusion.
@@
2. Related Work
2.1. Text-to-Image Diffusion Models
@@
Text-to-image diffusion models(Ramesh et al., 2022; Saharia et al., 2022; Nichol et al., 2021; Balaji et al., 2022) have achieved remarkable success in generating high-quality and diverse images. Early popular open-source models, such as Stable Diffusion(Rombach et al., 2022) and SDXL(Podell et al., 2023), are based on the U-Net(Ronneberger et al., 2015) architecture and utilize cross-attention between image and text to incorporate textual conditions. In contrast, Diffusion Transformer (DiT)(Peebles and Xie, 2023) replaces the U-Net backbone with a Transformer(Vaswani et al., 2017), demonstrating superior scalability and performance. Recent Multimodal-Diffusion Transformers (MM-DiTs), including Stable Diffusion 3(Esser et al., 2024) and FLUX.1(Labs, 2024), follow this design and project text and image tokens into a unified space to compute joint attention, thereby better integrating complex relationships between text and images. In this work, we implement our method on FLUX.1, where joint attention is computed as:
(1) $Q=\left[Q^{\text{txt}};Q^{\text{img}}\right],\quad K=\left[K^{\text{txt}};K^{\text{img}}\right],\quad V=\left[V^{\text{txt}};V^{\text{img}}\right]$
(2) $\tilde{Q}=PE\odot Q,\quad\tilde{K}=PE\odot K$
(3) $A=\operatorname{softmax}\left(\frac{\tilde{Q}\tilde{K}^{\top}}{\sqrt{d}}\right),\quad H_{\text{out}}=AV$
Here, $[;]$ denotes concatenation, $Q$, $K$, $V$ represent queries, keys, values, $PE$ is position embedding, $d$ denotes the dimension of the input features, $A$ denotes attention scores, and $H_{\text{out}}$ is the output of the attention layer.
2.2. Subject-Driven Generation
@@
Subject-driven generation(Ruiz et al., 2023; Gal et al., 2022; Kumari et al., 2023; Li et al., 2023; Ye et al., 2023) aims to synthesize novel scenes containing user-specified subjects. Early textual inversion methods(Gal et al., 2022; Dong et al., 2022; Voynov et al., 2023; Alaluf et al., 2023; Liu et al., 2023) learn new concepts by iteratively optimizing embeddings. Mainstream approaches fall into two categories: tuning-based and zero-shot. Tuning-based methods(Ruiz et al., 2023; Kumari et al., 2023; Gu et al., 2023) require per-subject fine-tuning to effectively capture visual details of the reference subject, which is resource-intensive and time-consuming, risks overfitting and catastrophic forgetting, and struggles with multi-subject composition. Zero-shot methods(Gal et al., 2023; Wei et al., 2023; Jia et al., 2023; Shi et al., 2024; Arar et al., 2023; Hua et al., 2023; Li et al., 2023; Ye et al., 2023; Ruiz et al., 2024; Wang et al., 2024a; Wu et al., 2025; Mou et al., 2025) introduce additional learnable modules to process extra image inputs, enabling personalization at inference time. Most of them rely on external encoders to compress images into embeddings, achieving only semantic-level alignment and suffering from insufficient visual consistency, particularly for novel concepts. UNO(Wu et al., 2025) and DreamO(Mou et al., 2025) leverage the VAE within base models to encode reference images for better subject consistency, yet they still struggle to retain fine-grained visual details (e.g., text and patterns). Furthermore, zero-shot methods require large-scale retraining when switching to new base models. Moreover, both categories can degrade the performance of base models, resulting in reduced quality of generated images.
@@
Beyond these categories, several training-free methods(Ding et al., 2024; Shin et al., 2024; Kang et al., 2025) aim to harness the in-context learning potential of pre-trained text-to-image models for plug-and-play subject-driven generation. FreeCustom(Ding et al., 2024) concatenates the reference images and noise in the latent space, and enhances the generated image’s focus on reference subjects by implementing a weighted mask strategy within the self-attention mechanism. DiptychPrompting(Shin et al., 2024) extends this paradigm to MM-DiT-based models via a ControlNet(Zhang et al., 2023), and reframes the subject-driven generation task as image completion on a diptych through meticulous prompt design. LatentUnfold(Kang et al., 2025) dispenses with ControlNet dependencies and evolves the diptych into a grid to enhance subject consistency. These methods avoid resource-demanding training or fine-tuning. However, constrained by the limited in-context learning capacity of base models, relying solely on self-attention struggles to fully reconstruct every detail of the reference subject. Our method identifies cross-image semantic correspondences while binding matched patches, reinforcing contextual connections between reference and generated subjects, thereby ensuring consistency across all components of the subject.
2.3. Appearance Transfer
Appearance transfer methods(Epstein et al., 2023; Alaluf et al., 2024; Mou et al., 2023, 2024; Go et al., 2024) aim to combine the appearance of one image with the structure of another. Techniques like Cross-Image (Alaluf et al., 2024) primarily rely on injecting key-value pairs across images to transfer appearance from reference to target positions in the generated image. Dragon-diffusion(Mou et al., 2023) and DiffEditor(Mou et al., 2024) integrate noise guidance with key-value injection. However, these methods often struggle to preserve structure accurately, as attention scores do not reliably reflect semantic correspondences, which can lead to erroneous matches. Methods like DIFT (Tang et al., 2023) address this by identifying fine-grained correspondences through pixel-pair similarity in U-Net’s feature space. Building on this, Eye-for-an-Eye(Go et al., 2024) achieves appearance transfer by directly replacing features of corresponding pixels. Nevertheless, these methods focus on style transfer rather than identity preservation, which can cause geometric inconsistencies (e.g., shape, limb proportions) between reference and generated subjects. Besides, these U-Net-based approaches are incompatible with modern MM-DiT architectures. In contrast, our work explores semantic matching and fine-grained feature transfer within the MM-DiT architecture while employing a dedicated noise initialization strategy to preserve the geometric structure of reference subjects.
Figure 3. Overview of FreeGraftor. First, we construct a collage based on the given text prompt and the reference image (Stage 1). Next, we invert this collage and record its diffusion trajectory (Stage 2). Finally, using the inverted noise as the initial latent representation, FreeGraftor synthesizes the output image through iterative denoising. During this process, the Semantic-aware Feature Grafting (SAFG) module integrates features from the collage to ensure alignment with the reference subject (Stage 3).
3. Method
3.1. Overview
@@
Our framework comprises three stages: collage construction, inversion, and generation (illustrated in Fig. 3). First, we construct a collage using the given reference image and text prompt. Next, we invert the collage to obtain initialization noise for the generation phase while extracting reference features. Finally, these features are injected into the generation process to synthesize the final image that contains the reference subject within the text-described scene.

### v4

@@ Abstract
Subject-driven image generation aims to synthesize novel scenes that faithfully preserve subject identity from reference images while adhering to textual guidance. However, existing methods struggle with a critical trade-off between fidelity and efficiency. Tuning-based approaches rely on time-consuming and resource-intensive subject-specific optimization, while zero-shot methods fail to maintain adequate subject consistency. In this work, we propose FreeGraftor, a training-free framework that addresses these limitations through cross-image feature grafting. Specifically, FreeGraftor leverages semantic matching and position-constrained attention fusion to transfer visual details from reference subjects to the generated images. Additionally, our framework introduces a novel noise initialization strategy to preserve the geometry priors of reference subjects, facilitating robust feature matching. Extensive qualitative and quantitative experiments demonstrate that our method enables precise subject identity transfer while maintaining text-aligned scene synthesis. Without requiring model fine-tuning or additional training, FreeGraftor significantly outperforms existing zero-shot and training-free approaches in both subject fidelity and text alignment. Furthermore, our framework can seamlessly extend to multi-subject generation, making it practical for real-world deployment. Our code is available at https://github.com/Nihukat/FreeGraftor.
I Introduction
With the rapid advancement of diffusion models[rombach2022high, podell2023sdxl, esser2024scaling, flux2024], text-to-image technology has evolved significantly, enabling photorealistic and diverse imagery generation from textual descriptions. Beyond generic synthesis, it has spawned personalized content creation tasks, among which subject-driven text-to-image generation is critical. This task aims to synthesize novel scenes with specific subjects via reference images and text prompts, requiring simultaneous high-fidelity subject preservation and precise text alignment.
Existing subject-driven methods face an inherent trade-off between fidelity and efficiency. These methods can be roughly divided into two categories based on whether test-time fine-tuning is required. One is tuning-based approaches (e.g., DreamBooth[ruiz2023dreambooth]), which adapt models through subject-specific fine-tuning on reference images. They achieve strong subject consistency but have notable drawbacks, including high computational cost, time consumption, catastrophic forgetting, and poor multi-subject handling. The other is zero-shot methods (e.g., IP-Adapter[ye2023ip]), which skip fine-tuning. Instead, they use auxiliary modules to extract and inject subject features from references. They offer high efficiency but fail to retain fine details, resulting in unsatisfactory alignment.
We identify that the breakthrough to this dilemma lies in the strong visual representation capabilities of pretrained text-to-image models. Specifically, cross-image semantic correspondences exist robustly in modern MM-DiT-based models (e.g., FLUX.1[flux2024]). As shown in Fig. 1, we compute token-level cosine similarities between the reference and target images in FLUX.1 during diffusion, and connect the most similar pixel pairs with colored lines. This visualization clearly shows that semantically similar regions also have high feature similarity in FLUX.1. This characteristic provides a reliable foundation for accurate subject feature transfer without extra training.
Figure 1: Cross-image semantic correspondence in FLUX.1. Semantically corresponding pixels between reference and generated images show high similarity in the model’s feature space.
@@ I Introduction
Leveraging this, we propose FreeGraftor, a training-free framework resolving the fidelity-efficiency dilemma via cross-image feature grafting. It extracts reference features through rectified flow inversion and injects them into generation dynamically. Specifically, we design a novel cross-image feature grafting mechanism transferring reference details to semantically matched regions, integrating semantic matching (for feature patch alignment) and position-constrained attention fusion (for reliable transfer). Additionally, we propose a structure-consistent initialization strategy which involves constructing a collage containing the subject and inverting it as generation initialization noise to ensure structural preservation. FreeGraftor achieves pixel-level detail retention with fast inference and low overhead, requiring no extra training.
@@
Extensive experiments demonstrate that our method outperforms existing approaches in both subject consistency and text alignment while maintaining competitive efficiency. Notably, it scales well to multi-subject scenarios, preserving individual identities and integrating subjects harmoniously into coherent scenes without layout conflicts or attribute confusion, expanding practical applications of subject-driven generation.
@@
II RELATED WORK
II-A Text-to-Image Diffusion Models
@@
Text-to-image diffusion models[rombach2022high, podell2023sdxl, esser2024scaling, flux2024] have achieved remarkable success in generating high-quality and diverse images. Early popular open-source models, such as Stable Diffusion[rombach2022high] and SDXL[podell2023sdxl], are based on the U-Net architecture and utilize cross-attention between image and text to incorporate textual conditions. In contrast, Diffusion Transformer (DiT)[peebles2023scalable] replaces the U-Net backbone with a Transformer, demonstrating superior scalability and performance. Recent Multimodal-Diffusion Transformers (MM-DiTs), including Stable Diffusion 3[esser2024scaling] and FLUX.1[flux2024], follow this design and project text and image tokens into a unified space to compute joint attention, thereby better integrating complex relationships between text and images.
II-B Subject-Driven Generation
@@
Subject-driven generation aims to synthesize novel scenes containing user-specified subjects. Mainstream approaches fall into two categories: tuning-based[ruiz2023dreambooth, kumari2023multi, gu2023mix] and zero-shot[li2023blip, ye2023ip, wu2025less, mou2025dreamo]. Tuning-based methods(e.g., Dreambooth[ruiz2023dreambooth]) require per-subject fine-tuning to effectively capture visual details of the reference subject, which is resource-intensive and time-consuming, risks overfitting and catastrophic forgetting, and struggles with multi-subject composition. Zero-shot methods introduce additional learnable modules to process extra image inputs, enabling personalization at inference time. Early methods (e.g., IP-Adapter[ye2023ip]) rely on ViT-based external encoders to compress images into embeddings, achieving only semantic-level alignment and suffering from insufficient visual consistency, particularly for novel concepts. Recent works (e.g., UNO[wu2025less]) leverage the VAEs within pretrained text-to-image models to encode reference images for better subject consistency, yet they still struggle to retain fine-grained visual details (e.g., text and patterns).
@@
Beyond these categories, several training-free methods[ding2024freecustom, shin2024large, kang2025flux] harness pre-trained models’ in-context learning ability for plug-and-play subject-driven generation. FreeCustom[ding2024freecustom] concatenates reference images and noise in latent space via self-attention weighted masking to focus on reference subjects. DiptychPrompting[shin2024large] extends to MM-DiT models via ControlNet[zhang2023adding], reframing the task as diptych image completion with tailored prompting. LatentUnfold[kang2025flux] dispenses with ControlNet and replaces diptych with grid for better consistency. These methods avoid heavy training but are limited by base models’ in-context learning capability, failing to fully reconstruct reference details via self-attention alone. Our method identifies cross-image semantic correspondences, binds matched patches, and reinforces reference-generated contextual connections to ensure full component consistency.
III METHOD
III-A Overview
@@
Our framework comprises three stages: collage construction, inversion, and generation (Fig. 2). First, we construct a collage using the given reference image and text prompt. Next, we invert the collage to obtain initialization noise for the generation phase while extracting reference features. Finally, these features are injected into the generation process to synthesize the final image that contains the reference subject within the text-described scene.
Figure 2: Overview of FreeGraftor. The framework consists of three stages: (1) Collage construction using reference image and text prompt, (2) Inversion to obtain initialization noise, and (3) Generation with feature grafting.
III-B Semantic-Aware Feature Grafting
Inspired by appearance transfer techniques[tang2023emergent, go2024eye], we leverage the inherent capabilities of text-to-image diffusion models to transfer visual characteristics of reference subjects to corresponding regions in generated images without incurring additional training costs. To achieve this, we propose the Semantic-Aware Feature Grafting (SAFG) module, which first establishes semantic correspondences between reference and generated patches via semantic matching, then fuses features through position-constrained attention fusion (Fig. 3).
III-B1 Semantic Matching
Early appearance transfer methods directly concatenate keys and values of reference and generated images to compute cross-image attention. However, high attention scores do not necessarily indicate semantic correspondences across images, and erroneous matches may lead to structural distortions and visual artifacts. DIFT[tang2023emergent] revealed that U-Net-based diffusion models contain rich semantic information in their feature space and established cross-image correspondences via cosine similarity-based semantic matching. We extend this principle to MM-DiT-based models.
Given a reference image $I^{\text{ref}}$ and a predefined binary mask $M^{\text{ref, pre}}$ indicating the reference subject, we extract attention features using rectified flow inversion and match reference $H^{\text{ref}}$ and generated $H^{\text{gen}}$ features at the same timestep and attention layer. For each reference patch $i$ within $M^{\text{ref, pre}}$, we compute cosine similarity with all generated patches $j$:
$s_{ij}=\frac{H_{i}^{\text{img,ref}}\cdot H_{j}^{\text{img,gen}}}{\|H_{i}^{\text{img,ref}}\|_{2}\|H_{j}^{\text{img,gen}}\|_{2}},\quad m(i)=\arg\max_{j}s_{ij}$ (1)
$M_{i}^{\text{ref, sim}}=\begin{cases}1,&s_{i,m(i)}\geq\tau\\ 0,&s_{i,m(i)}<\tau\end{cases}$ (2)
ii. Cycle Consistency Filtering. Following DreamMatcher[nam2024dreammatcher], for each reference patch $i$ (satisfying $M_{i}^{\text{ref, pre}}=1$), we find the most similar reference patch $k$ to its matched generated patch $m(i)$:
$m^{-1}(m(i))=\arg\max_{k:M_{k}^{\text{ref,pre}}=1}\frac{H_{m(i)}^{\text{img,gen}}\cdot H_{k}^{\text{img,ref}}}{\|H_{m(i)}^{\text{img,gen}}\|_{2}\|H_{k}^{\text{img,ref}}\|_{2}}$ (3)
$d_{i}=\|\mathbf{p}_{i}^{\text{ref}}-\mathbf{p}_{m^{-1}(m(i))}^{\text{ref}}\|_{2},\quad M_{i}^{\text{ref, consi}}=\begin{cases}1,&d_{i}\leq\delta\\ 0,&d_{i}>\delta\end{cases}$ (4)
$M^{\text{ref}}=M^{\text{ref, pre}}\odot M^{\text{ref, sim}}\odot M^{\text{ref, consi}}$ (5)
III-B2 Position-Constrained Attention Fusion
$K^{\text{cat}}=[K^{\text{txt}};K^{\text{img,gen}};K_{M}^{\text{img,ref}}],$ (6)
$V^{\text{cat}}=[V^{\text{txt}};V^{\text{img,gen}};V_{M}^{\text{img,ref}}]$
$PE^{\text{cat}}=[PE^{\text{txt}};PE^{\text{img,gen}};\{PE_{m(i)}^{\text{img,gen}}\mid M_{i}^{\text{ref}}=1\}]$ (7)
The revised attention computation is formulated as:
$\tilde{Q}=PE\odot Q,\quad\tilde{K}^{\text{cat}}=PE^{\text{cat}}\odot K^{\text{cat}}$ (8)
$A^{+}=\text{softmax}\left(\frac{\tilde{Q}(\tilde{K}^{\text{cat}})^{\top}}{\sqrt{d}}\right),\quad H_{\text{out}}^{+}=A^{+}V^{\text{cat}}$ (9)
III-C Structure-Consistent Initialization
To prevent the generated subject from overfitting the pose of the reference subject, we introduce a dynamic feature dropout strategy. Since structural layouts are predominantly determined during early diffusion steps, we reduce the injection of reference features during the early steps while retaining them in the later steps to preserve visual details. At timestep $t$, for each reference patch, we apply dropout with probability $p=\omega t$, where $\omega\in[0,1]$ controls dropout intensity. The dropout mask is represented by a Bernoulli distribution:
$M^{\text{drop}}_{i,j}\sim\text{Bernoulli}(1-\omega t)\quad\forall i,j$ (10)
The final reference mask is computed by applying dropout to the reference image mask obtained from semantic matching:
${M^{\text{ref}}}^{\prime}=M^{\text{ref}}\odot M^{\text{drop}}$ (11)
This strategy preserves the reference subject’s structure while aligning it with the text-driven layout and maintaining pose flexibility. For multi-subject generation, we iteratively replace subjects in the template and invert the multi-subject collage, avoiding simultaneous processing of multiple reference images and eliminating computational overhead.
Figure 3: Semantic-Aware Feature Grafting (SAFG) Module. The module establishes semantic correspondence and fuses features through position-constrained attention.
IV EXPERIMENTS
TABLE I: Comparing to other subject-driven generation methods, our method achieves the best performance.
Image-alignment Text-alignment
[…]
FreeCustom SD1.5 0.8308 0.6107 0.3246 0.6223
BLIP-Diffusion SD1.5 0.8621 0.7364 0.2949 0.2360
MS-Diffusion SDXL 0.9023 0.7977 0.3254 1.3405
IP-Adapter FLUX 0.8920 0.7696 0.3048 0.7444
UNO FLUX 0.9069 0.8237 0.3151 1.2788
DreamO FLUX 0.9159 0.8400 0.3217 1.6234
DiptychPrompting FLUX 0.8924 0.7971 0.3291 1.5728
LatentUnfold FLUX 0.8741 0.7755 0.3170 1.3929
Ours FLUX $\mathbf{0.9527}$ $\mathbf{0.9042}$ $\mathbf{0.3308}$ $\mathbf{1.6481}$
Figure 4: Single-subject generation results. FreeGraftor preserves pixel-level details while enabling flexible text-guided control.
Figure 5: Ablation study results. Each component of FreeGraftor is critical for achieving high-quality subject-driven generation.
IV-A Experimental Setup
IV-A1 Implementation Details
Our method is implemented on FLUX.1-dev. For collage construction, we use Grounding DINO[liu2024grounding] to localize subjects, SAM[kirillov2023segment] to segment subjects, and LaMa[suvorov2022resolution] to erase the original subject in the template image. FireFlow[deng2024fireflow] is employed for rectified flow inversion. Template generation, inversion, and final generation each take 25 diffusion steps. During semantic matching, we set the similarity threshold $\tau=0.2$, cycle consistency threshold $\delta=1.5$, and dropout intensity $\omega=0.5$.
IV-A2 Baselines
We compare our method with zero-shot methods (BLIP-Diffusion[li2023blip], IP-Adapter[ye2023ip], MS-Diffusion[wang2024ms], UNO[wu2025less], DreamO[mou2025dreamo]) and training-free methods (FreeCustom[ding2024freecustom], DiptychPrompting[shin2024large], LatentUnfold[kang2025flux]). Details of these methods are provided in Appendix B.
IV-A3 Datasets and Evaluation
For qualitative evaluation, we use reference images from DreamBench[ruiz2023dreambooth], CustomConcept101[kumari2023multi], and Mix-of-Show[gu2023mix] for subject-driven generation. For quantitative evaluation, we generate 30×25×4=3,000 images using 30 subjects and 25 prompt templates from DreamBench[ruiz2023dreambooth], with 4 random seeds per subject-prompt pair. We evaluate performance from two perspectives: text alignment and image alignment. For text alignment, we report results using CLIP[radford2021learning] and ImageReward[xu2024imagereward] metrics. For image alignment, we use GroundedSAM[ren2024grounded] to extract segments of reference and generated subjects, respectively, and compute similarity scores between these segments using CLIP[radford2021learning] and DINOv2[oquab2023dinov2].
TABLE II: Comparing to ablation variants, our full method achieves the best performance.
Image-alignment Text-alignment
[…]
w/o Init 0.8529 0.7138 0.3236 1.6320
w/o Graft 0.8463 0.6572 0.3306 1.6434
w/o Match 0.9504 0.8974 0.3232 1.5258
w/ replace 0.8476 0.7010 0.3174 1.4551
Ours $\mathbf{0.9527}$ $\mathbf{0.9042}$ $\mathbf{0.3308}$ $\mathbf{1.6481}$
IV-B Quantitative Results
As shown in Table I, our method achieves the highest alignment across all metrics, demonstrating superior subject fidelity and textual adherence, validating its effectiveness for subject-driven generation.
IV-C Qualitative Results
Fig. 4 presents the results for single-subject generation using FreeGraftor and other methods. FreeCustom, BLIP-Diffusion and IP-Adapter fail to retain fine visual details of the reference subject effectively. While other baselines partially recover reference characteristics, they struggle with fine-grained fidelity. FreeGraftor achieves pixel-level detail preservation, such as logos on the can, numbers on the clock, and text on the teddy bear. For non-rigid subjects (e.g., the teddy bear), FreeGraftor generates natural pose variations aligned with text guidance, demonstrating the flexibility and robustness of our feature grafting technique.
IV-D Ablation Study
To validate each design’s effectiveness in our method, we compare several ablation variants in Figure 5. (a) When structure-consistent initialization is removed, the generated images inherit the reference subject’s appearance (e.g., colors, patterns) but fail to maintain structural consistency (e.g., distorted body proportions in dogs). (b) Without feature grafting, the generated subjects lack identifiable correspondence with the reference subjects. (c) If semantic matching is omitted and reference keys and values are directly appended during attention computation, the generated subjects exhibit rigid poses and artifacts. (d) Replacing position-constrained attention fusion with direct feature replacement introduces incoherent edges and distortions. Quantitative results in Table II confirm that both structure-consistent initialization and feature grafting significantly improve image alignment, as they are critical for preserving subject identity. Skipping semantic matching or positional fusion degrades text alignment due to reduced image quality.
V Conclusion
We present FreeGraftor, a novel training-free framework for subject-driven generation that bridges the critical gap between subject fidelity and computational efficiency. By leveraging Semantic-Aware Feature Grafting, FreeGraftor seamlessly injects the appearance of reference subjects into generated scenes. Additionally, our Structure-Consistent Initialization strategy preserves reference geometry and enhances the robustness of semantic matching. Extensive experiments demonstrate that FreeGraftor achieves pixel-level detail preservation and flexible text-guided control while requiring no training or test-time optimization. Our framework can be seamlessly extended to multi-subject generation, faithfully retaining visual details of all reference subjects while incurring no additional computational overhead.
-A Pseudo Code
.0-A1 FreeGraftor Pipeline
The complete FreeGraftor pipeline is outlined in Algorithm 1, detailing the three main stages: collage construction, inversion, and generation.